# Preference-Conditioned Gradient Variations for Multi-Objective Quality-Diversity

### v1

@@ II-C Multi-Objective Quality-Diversity Algorithms
Multi-Objective Quality-Diversity (moqd) combines the goals of QD and MO optimisation. Specifically, the goal of moqd is to return the Pareto front of solutions in each cell of the feature space with maximum hypervolume, $\mathcal{P}(\mathcal{C}_{i})$ [21]. This moqd goal can be mathematically formulated as:
$\max_{\theta\in\Theta}\sum_{i=1}^{k}\Xi(\mathcal{P}_{i}),\,\,\text{where}\,\,\forall i,\mathcal{P}_{i}=\mathcal{P}({\theta|\Phi(\theta)\in\mathcal{C}_{i}})$ (4)
@@
mome [21] was the first moqd algorithm that aimed to achieve this moqd goal. To achieve this, mome maintains a Pareto front in each cell of a MAP-Elites grid. At each iteration, a cell is uniformly selected and then a solution from the corresponding Pareto front is uniformly selected. Then, the algorithm follows a standard map-elites procedure: the solution undergoes genetic variation and is evaluated. The evaluated solution is added back to the grid if it lies on the Pareto front of the cell corresponding to its feature vector.
II-D Problem Formulation
In this work, we consider an agent sequentially interacting with an environment for an episode of length $T$, modelled as a Multi-Objective Markov Decision Process (momdp), defined by $\langle\mathcal{S},\mathcal{A},\mathcal{P},\textbf{\em R},\Omega\rangle$. At each discrete time step $t$, the agent observes the current state $s_{t}\in\mathcal{S}$ and takes an action $a_{t}\in\mathcal{A}$ by following a policy $\pi_{\theta}$ parameterized by $\theta$. Consequently, the agent transitions to a new state sampled from the dynamics probability distribution $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})$. The agent also receives a reward vector $\mathbf{r}_{t}=[r_{1}(s_{t},a_{t}),\dots,r_{m}(s_{t},a_{t})]$, where each reward function $r_{i}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ defines an objective. The multi-objective fitness of a policy $\pi$ is defined as a vector F$(\pi)=[f_{1}(\pi),...,f_{m}(\pi)]$. Here, each $f_{i}$ represents the expected discounted sum of rewards, calculated as $f_{i}=\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r_{i}(s_{t},a_{t})\right]$ for a given reward function $r_{i}$. The discount rate $\gamma\in[0,1]$ controls the relative weighting of immediate and long-term rewards.
II-E Reinforcement Learning
@@
In the single-objective case ($m=1$), the momdp collapses into a simple Markov Decision Process (MDP) with scalar rewards, where the goal is to find a policy $\pi$ that maximises the expected discounted sum of rewards or return, $F(\pi)=\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t})\right]$. Numerous Reinforcement Learning (rl) methods have been developed to address the challenge of finding policies that optimize this cumulative reward. One particularly relevant approach is the Twin Delayed Deep Deterministic Policy Gradient algorithm (td3)[34].
@@
The td3 algorithm belongs to the broader family of actor-critic rl techniques [35], which involve two key components: an actor network and a critic network. The actor network is a policy parameterised by $\phi$, denoted $\pi_{\phi}$ that is used to interact with the environment. The transitions $(s_{t},a_{t},r_{t},s_{t+1})$ coming from the interactions with the environment are stored in a replay buffer $\mathcal{B}$ and used to train the actor and the critic. The critic network is an action-value function parameterised by $\psi$, denoted $Q_{\psi}$ that evaluates the quality of the actor’s actions and helps the agent learn to improve its decisions over time. The critic estimates the expected return obtained when starting from state $s$, taking action $a$ and following policy $\pi$ thereafter, $Q_{\psi}(s,a)=\mathbb{E}_{\pi}[\sum_{t}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
@@
$y=r(s_{t},a_{t})+\gamma\min_{i=1,2}Q_{\psi_{i}}(s_{t+1},\pi_{\phi^{\prime}}(s_{t+1})+\epsilon)$ (5)
@@
$\nabla_{\phi}J(\pi_{\phi})=\mathbb{E}\big{[}\,\nabla_{\phi}\,\pi_{\phi}(s)\,\nabla_{a}Q_{\psi_{1}}(s,a)\,|\,_{a=\pi_{\phi}(s)}\big{]}$ (6)
@@
III Related Works
III-A Multi-Objective Evolutionary Algorithms
Multi-Objective Evolutionary Algorithms (moea) evolve a population of potential solutions iteratively over several generations to identify an optimal set of solutions that balance conflicting objectives. At each iteration, solutions are selected from the population and undergo genetic variation (through crossover and mutation operators) and are then added back to the population. Different MOEAs can vary in terms of their specific selection strategies, crossover and mutation operators, population management techniques, and how they maintain diversity in the population [36].
Non-dominated Sorting Genetic Algorithm II nsga-ii [19] and Strength Pareto Evolutionary Algorithm 2 (spea2) [18] both use biased selection mechanisms to guide the optimisation process. Both methods select solutions that are higher performing and occupy less dense regions of the objective space with higher probability. This guides the population towards higher-performing Pareto fronts, while simultaneously ensuring solutions are well-distributed across the front.
Our method, mome-p2c has synergies with many methods from moea literature including biased selection and addition mechanisms (see Section IV-A) and we refer the interested reading to a comprehensive survey of moea algorithms for more details [36]. However, our method differs from traditional moea approaches in two significant aspects. First, it employs a map-elites grid to explicitly maintain solutions that are diverse in feature space while optimising over objectives. Second, it incorporates techniques from reinforcement learning to form gradient-based mutations which help to overcome the limited search power of traditional ga variations for high-dimensional search spaces [22].
III-B Multi-Objective Reinforcement Learning
@@
$\begin{split}J(\pi,\bm{\omega})=\mathbb{E}_{\pi}\Big{[}\sum_{t}\bm{\omega}^{\intercal}\bm{r}_{t}\Big{]}=\bm{\omega}^{\intercal}\mathbb{E}_{\pi}\Big{[}\sum_{t}\bm{r}_{t}\Big{]}=\bm{\omega}^{\intercal}\bm{J}(\pi)\end{split}$ (7)
@@
Here, $\bm{\omega}$ is referred to as a preference, with $\sum_{i}\omega_{i}=1$. The preference quantifies the relative importance of each of the objective functions for the end-user and, when the preference is fixed, we can collapse the momdp into a single-objective setting that can be optimised with well-established rl approaches.
@@
In single-policy approaches, the challenge arises in determining the preference vector beforehand, as it may prove to be a complex task or may vary among different users [32]. Instead, it may be useful to find solutions which are optimal for different preference values so that the user can examine the range of possible solutions that is on offer and then assign their preferences retrospectively [32]. With this perspective in mind, multi-policy morl methods aim to find a set of policies that excel across a range of different preferences [37, 38]. Often, each policy in the set is trained using preference-conditioned policy-gradient derived from a multi-objective, preference-conditioned action-value function [37, 38, 39].
@@
Some methods straddle the line between single-policy and multi-policy morl by seeking a single preference-conditioned policy that can maximise the weighted sum of expected returns (Equation 7) for any given preference [40, 41, 39, 42]. This approach offers advantages such as reduced storage costs and rapid adaptability [41]. However, while having preference-conditioned policy approaches might be cheaper and more flexible, these methods have been observed to achieve worse performance on the objective functions for any given preference than having specialised policies [38].
@@
Our work combines elements of both preference-conditioned and multi-policy approaches. Our actor-critic networks are preference-conditioned. However, within each cell of the map-elites grid, we adopt a multi-policy approach. While storing many policies in each cell is more costly in terms of memory, relying solely on a single preference-conditioned policy in each grid cell would not offer a straightforward means to assess whether a new solution is superior or not. One possible strategy would be to evaluate each policy on a predefined set of preferences, and replace the policy in the grid if it achieves higher values on those preferences. However, this would require multiple costly evaluations so this approach is not practical. To the best of our knowledge, there is no prior research in multi-objective reinforcement learning (morl) that actively seeks to diversify behaviours in this manner.
III-C Gradients in Quality-Diversity
@@
qd algorithms belong to the wider class of Genetic Algorithms (ga), which broadly adhere to a common structure of selection, variation and addition to a population of solutions. While these methods have been observed to be highly-effective black box methods, one key limitation is their lack of scalability to high-dimensional search spaces. In tasks in which solutions are the parameters of a neural network, the search space can be thousands of dimensions and thus traditional ga variation operators do not provide sufficient exploration power. To address this, many works in single-objective qd leverage the search power of gradient-based methods in high-dimensional search spaces [22, 23, 24, 43, 44, 45]. The pioneer of these methods, Policy-gradient assisted map-elites (pga-me) [22], combines the td3 algorithm with the map-elites algorithms to apply qd to high-dimensional robotics control tasks. In particular, during the evaluation of solutions in pga-me, environment transitions are stored and used to train actor and critic networks, using the training procedures explained in Section II-E. Then, pga-me follows a normal map-elites loop except in each iteration, half of the solutions are mutated via ga variations and the other half are mutated via policy-gradient (pg) updates.
@@
Since pga-me, several other qd algorithms with gradient-based variation operators have been proposed. Some of these are tailored to consider different task settings which have differentiable objective and feature functions [43] or discrete action spaces [46]. Other methods use policy gradient updates to improve both the fitness and diversity of solutions [24, 43, 44]. A particular method of note is dcg-me [45, 47] which uses policy-gradient variations conditioned on features of solutions. Similar to mome-p2c, the motivation for this method is to provide more nuanced gradient information. Conditioning the policy-gradient on the feature value of a solution provides a way to update the solution toward higher performance, given that it has a certain behaviour. However, this method only considered mono-objective problems. Other than mome-pgx (see Section III-D) we are unaware of gradient-based qd methods applied multi-objective problems.
III-D Multi-Objective Quality-Diversity Algorithms
@@
Recently, policy gradient variations, inspired by single-objective methods, have played a pivotal role in shaping the development of techniques in moqd. Notably, while mome (see Section III-D) is a simple and effective moqd approach, it relies on ga policy-gradient mutations as an exploration mechanism which makes it inefficient in high-dimensional search spaces. To overcome this challenge, Multi-Objective MAP-Elites with Policy-Gradient Assistance and Crowding-based Exploration (mome-pgx) [25] was recently introduced as an effort to improve the performance and data-efficiency of mome in tasks that can be framed as a momdp. mome-pgx maintains an actor and critic network for each objective function separately and uses policy gradient mutation operators in order to drive better exploration in the solution search space. mome-pgx also uses crowding-based selection and addition mechanisms to bias exploration in sparse regions of the Pareto front and to maintain a uniform distribution of solutions on the front. mome-pgx was shown to outperform mome and other baselines across a suite of multi-objective robotics tasks involving high-dimensional neural network policies. Despite this success, mome-pgx requires maintaining distinct actor-critic pairs for each objective, which is costly in memory. Moreover, since each actor-critic network pair learns about each of the objective separately, the pg variations may only provide disjoint gradient information about each of the objectives, and fail to capture nuanced trade-offs.
@@
To the best of our knowledge mome and mome-pgx are the only existing moqd algorithms to date. However, we also note of two particularly relevant approaches which have synergies with the moqd setting. Multi-Criteria Exploration (mcx) [12] which uses a tournament ranking strategy to condense a solution’s score across multiple objectives into a single value, and then uses a standard MAP-Elites strategy. Similarly, Many-objective Optimisation via Voting for Elites (move) [26] uses a MAP-Elites grid to find solutions which are high-performing on many-objective problems. In this method, each cell of the grid represents a different subset of objectives and a solution replaces the existing solution in the cell if it is better on at least half of the objectives for the cell. While both mcx and move consider the simultaneous maximisation of many objectives, they both aim to find one solution per cell in the map-elites grid rather than Pareto fronts for different features. Therefore, we consider their goals to be fundamentally different from the moqd goal defined in Equation 4.
IV MOME-P2C

### v2

@@ 2.3. Multi-Objective Quality-Diversity Algorithms
Multi-Objective Quality-Diversity (moqd) combines the goals of QD and MO optimisation. Specifically, the goal of moqd is to return the non-dominated front of solutions in each cell of the feature space with maximum hypervolume, $\mathcal{P}(\mathcal{C}_{i})$ (Pierrot et al., 2022b). This moqd goal can be mathematically formulated as:
(4) $\max_{\theta\in\Theta}\sum_{i=1}^{k}\Xi(\mathcal{P}_{i}),\,\,\text{where}\,\,\forall i,\mathcal{P}_{i}=\mathcal{P}({\theta|\Phi(\theta)\in\mathcal{C}_{i}})$
@@
mome (Pierrot et al., 2022b) was the first moqd algorithm that aimed to achieve this moqd goal. To achieve this, mome maintains a non-dominated front in each cell of a MAP-Elites grid. At each iteration, a cell is uniformly selected and then a solution from the corresponding non-dominated front is uniformly selected. Then, the algorithm follows a standard map-elites procedure: the solution undergoes genetic variation and is evaluated. The evaluated solution is added back to the grid if it lies on the non-dominated front of the cell corresponding to its feature vector.
We note that moqd differs from related areas such as multi-modal multi-objective optimisation (Li et al., 2023; Liang et al., 2016). The purpose of multi-modal multi-objective optimisation is to provide a set of equivalent Pareto-optimal outcomes, where each outcome achieves nearly the same balance of objectives but corresponds to a different solution representation, giving end-users flexibility in practice. For example, finding two different circuit layouts that yield the same power–efficiency trade-off, yet offer engineers different implementation choices. In other words, in multi-modal multi-objective optimisation, the goal is to identify multiple Pareto-optimal solutions that are equivalent in objective space but differ in decision space. By contrast, moqd explicitly introduces a feature space $\Phi(\theta)$ that structures the search according to user-defined features. The aim is then to return a Pareto front approximation within each feature cell, rather than globally across all solutions. This distinction means that while multi-modal multi-objective optimisation emphasises diversity of solutions that achieve similar trade-offs, moqd emphasises diversity across features and simultaneously seeks high-performing trade-offs within each feature category.
2.4. Problem Formulation
[…]
2.5. Reinforcement Learning
@@
In the single-objective case ($m=1$), the momdp collapses into a simple Markov Decision Process (MDP) with scalar rewards, where the goal is to find a policy $\pi$ that maximises the expected discounted sum of rewards or return, $F(\pi)=\mathbb{E}_{\pi}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t})\right]$. Numerous Reinforcement Learning (rl) methods have been developed to address the challenge of finding policies that optimize this cumulative reward. One particularly relevant approach is the Twin Delayed Deep Deterministic Policy Gradient algorithm (td3)(Fujimoto et al., 2018).
@@
The td3 algorithm belongs to the broader family of actor-critic rl techniques (Grondman et al., 2012), which involve two key components: an actor network and a critic network. The actor network is a policy parameterised by $\phi$, denoted $\pi_{\phi}$ that is used to interact with the environment. The transitions $(s_{t},a_{t},r_{t},s_{t+1})$ coming from the interactions with the environment are stored in a replay buffer $\mathcal{B}$ and used to train the actor and the critic. The critic network is an action-value function parameterised by $\psi$, denoted $Q_{\psi}$ that evaluates the quality of the actor’s actions and helps the agent learn to improve its decisions over time. The critic estimates the expected return obtained when starting from state $s$, taking action $a$ and following policy $\pi$ thereafter, $Q_{\psi}(s,a)=\mathbb{E}_{\pi}[\sum_{t}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
@@
(5) $y=r(s_{t},a_{t})+\gamma\min_{i=1,2}Q_{\psi_{i}}(s_{t+1},\pi_{\phi^{\prime}}(s_{t+1})+\epsilon)$
@@
(6) $\nabla_{\phi}J(\pi_{\phi})=\mathbb{E}\big[\,\nabla_{\phi}\,\pi_{\phi}(s)\,\nabla_{a}Q_{\psi_{1}}(s,a)\,|\,_{a=\pi_{\phi}(s)}\big]$
@@
3. Related Works
3.1. Multi-Objective Evolutionary Algorithms
Multi-Objective Evolutionary Algorithms (moea) evolve a population of potential solutions iteratively over several generations to identify an optimal set of solutions that balance conflicting objectives. At each iteration, solutions are selected from the population and undergo genetic variation (through crossover and mutation operators) and are then added back to the population. Different MOEAs can vary in terms of their specific selection strategies, crossover and mutation operators, population management techniques, and how they maintain diversity in the population (Zhou et al., 2011).
Non-dominated Sorting Genetic Algorithm II (nsga-ii) (Deb et al., 2000) and Strength Pareto Evolutionary Algorithm 2 (spea2) (Zitzler et al., 2001) both use non-uniform selection mechanisms to guide the optimisation process. Both methods select solutions that are higher performing and occupy less dense regions of the objective space with higher probability. This guides the population towards higher-performing fronts, while simultaneously ensuring solutions are well-distributed across the front.
Our method, mome-p2c has synergies with many methods from moea literature including non-uniform selection and addition mechanisms (see Section 4.1) and we refer the interested reading to a comprehensive survey of moea algorithms for more details (Zhou et al., 2011). However, we emphasise that moea algorithms treat all objectives as quantities to maximise, seeking trade-offs among them. However, in moqd algorithms, the goal is not to maximise feature values but to cover the feature space with high-performing solutions. In other words, moqd algorithms seek solutions which are diverse in the feature-space, and then to maximise objectives for each feature. Moreover, our method differs from traditional moea approaches in two significant aspects. First, it employs a map-elites grid to explicitly maintain solutions that are diverse in feature space while optimising over objectives. Second, it incorporates techniques from reinforcement learning to form gradient-based mutations which help to overcome the limited search power of traditional ga variations for high-dimensional search spaces (Nilsson and Cully, 2021).
3.2. Multi-Objective Reinforcement Learning
@@
(7) $\begin{split}J(\pi,\bm{\omega})=\mathbb{E}_{\pi}\Big[\sum_{t}\bm{\omega}^{\intercal}\bm{r}_{t}\Big]=\bm{\omega}^{\intercal}\mathbb{E}_{\pi}\Big[\sum_{t}\bm{r}_{t}\Big]=\bm{\omega}^{\intercal}\bm{J}(\pi)\end{split}$
@@
Here, $\bm{\omega}$ is referred to as a preference, with $\sum_{i}\omega_{i}=1$. The preference quantifies the relative importance of each of the objective functions and, when the preference is fixed, we can collapse the momdp into a single-objective setting that can be optimised with well-established rl approaches.
@@
In single-policy approaches, the challenge arises in determining the preference vector beforehand, as it may prove to be a complex task or may vary among different users (Hayes et al., 2022). Instead, it may be useful to find solutions which are optimal for different preference values so that the user can examine the range of possible solutions that is on offer and then assign their preferences retrospectively (Hayes et al., 2022). With this perspective in mind, multi-policy morl methods aim to find a set of policies that excel across a range of different preferences (Mossalam et al., 2016; Xu et al., 2020). Often, each policy in the set is trained using preference-conditioned policy-gradient derived from a multi-objective, preference-conditioned action-value function (Mossalam et al., 2016; Xu et al., 2020; Yang et al., 2019).
@@
Some methods straddle the line between single-policy and multi-policy morl by seeking a single preference-conditioned policy that can maximise the weighted sum of expected returns (Equation 7) for any given preference (Parisi et al., 2016; Abels et al., 2019; Yang et al., 2019; Zhu et al., 2023). This approach offers advantages such as reduced storage costs and rapid adaptability (Abels et al., 2019). However, while having preference-conditioned policy approaches might be cheaper and more flexible, these methods have been observed to achieve worse performance on the objective functions for any given preference than having specialised policies (Xu et al., 2020).
@@
Our work combines elements of both preference-conditioned and multi-policy approaches. Our actor-critic networks are preference-conditioned and are separate from the map-elites grid. Hence this preference-conditioned actor can be seen as a generalist policy. However, within each cell of the map-elites grid, we store specialised, non-conditioned policies hence taking a multi-policy approach. To the best of our knowledge, there is no prior research in multi-objective reinforcement learning (morl) that actively seeks to diversify behaviours in this manner.
3.3. Gradients in Quality-Diversity
@@
qd algorithms belong to the wider class of Genetic Algorithms (ga), which broadly adhere to a common structure of selection, variation and addition to a population of solutions. While these methods have been observed to be highly-effective black box methods, one key limitation is their lack of scalability to high-dimensional search spaces. In tasks in which solutions are the parameters of a neural network, the search space can be thousands of dimensions and thus traditional ga variation operators do not provide sufficient exploration power. To address this, many works in single-objective qd leverage the search power of gradient-based methods in high-dimensional search spaces (Nilsson and Cully, 2021; Flageat et al., 2022; Pierrot et al., 2022a; Fontaine and Nikolaidis, 2021; Tjanaka et al., 2022; Faldor et al., 2023). The pioneer of these methods, Policy-gradient assisted map-elites (pga-me) (Nilsson and Cully, 2021), combines the td3 algorithm with the map-elites algorithms to apply qd to high-dimensional robotics control tasks. In particular, during the evaluation of solutions in pga-me, environment transitions are stored and used to train actor and critic networks, using the training procedures explained in Section 2.5. Then, pga-me follows a normal map-elites loop except in each iteration, half of the solutions are mutated via ga variations and the other half are mutated via policy-gradient (pg) updates.
@@
Since pga-me, several other qd algorithms with gradient-based variation operators have been proposed. Some of these are tailored to consider different task settings which have differentiable objective and feature functions (Fontaine and Nikolaidis, 2021) or discrete action spaces (Boige et al., 2023). Other methods use policy gradient updates to improve both the fitness and diversity of solutions (Pierrot et al., 2022a; Fontaine and Nikolaidis, 2021; Tjanaka et al., 2022). A particular method of note is dcg-me (Faldor et al., 2023, 2024) which uses policy-gradient variations conditioned on features of solutions. Similar to mome-p2c, the motivation for this method is to provide more nuanced gradient information. Conditioning the policy-gradient on the feature value of a solution provides a way to update the solution toward higher performance, given that it has a certain behaviour. However, this method only considered mono-objective problems. Other than mome-pgx (see Section 3.4) we are unaware of gradient-based qd methods applied multi-objective problems.
3.4. Multi-Objective Quality-Diversity Algorithms
@@
Recently, policy gradient variations, inspired by single-objective methods, have played a pivotal role in shaping the development of techniques in moqd. Notably, while mome (see Section 3.4) is a simple and effective moqd approach, it relies on ga policy-gradient mutations as an exploration mechanism which makes it inefficient in high-dimensional search spaces. To overcome this challenge, Multi-Objective MAP-Elites with Policy-Gradient Assistance and Crowding-based Exploration (mome-pgx) (Janmohamed et al., 2023) was recently introduced as an effort to improve the performance and data-efficiency of mome in tasks that can be framed as a momdp. mome-pgx maintains an actor and critic network for each objective function separately and uses policy gradient mutation operators in order to drive better exploration in the solution search space. mome-pgx also uses crowding-based selection and addition mechanisms to bias exploration in sparse regions of the non-dominated front and to maintain a uniform distribution of solutions on the front. mome-pgx was shown to outperform mome and other baselines across a suite of multi-objective robotics tasks involving high-dimensional neural network policies. However, since each actor-critic network pair learns about each of the objective separately, the pg variations in mome-pgx may only provide disjoint gradient information about each of the objectives, and fail to capture nuanced trade-offs.
@@
To the best of our knowledge mome and mome-pgx are the only existing moqd algorithms to date. However, we also note of two particularly relevant approaches which have synergies with the moqd setting. Multi-Criteria Exploration (mcx) (Gaier et al., 2022) which uses a tournament ranking strategy to condense a solution’s score across multiple objectives into a single value, and then uses a standard MAP-Elites strategy. Similarly, Many-objective Optimisation via Voting for Elites (move) (Dean and Cheney, 2023) uses a MAP-Elites grid to find solutions which are high-performing on many-objective problems. In this method, each cell of the grid represents a different subset of objectives and a solution replaces the existing solution in the cell if it is better on at least half of the objectives for the cell. While both mcx and move consider the simultaneous maximisation of many objectives, they both aim to find one solution per cell in the map-elites grid rather than non-dominated fronts for different features. Therefore, we consider their goals to be fundamentally different from the moqd goal defined in Equation 4.
Additionally, several works have used multi-objective mechanisms to enhance single-objective qd. These works consider quality (fitness) and diversity directly as optimisation objectives, rather than task-based objectives. For instance, novelty search with local competition (Lehman and Stanley, 2011) balances novelty and performance through a multi-objective selection scheme, indicator-based methods (Neumann et al., 2019) employ measures such as hypervolume to guide the spread of solutions, and Pareto-based selection approaches (Wang et al., 2023; Do et al., 2024) explicitly use diversity criteria in archive updates. In all cases, diversity acts as a meta-objective that shapes the search, but the final output remains a single repertoire of diverse, high-quality solutions, in contrast to moqd which seeks Pareto fronts of task-based trade-offs within each feature.
4. MOME-P2C